The largest meteorite in the world is hoba west located in Namibia. It's weight is three times more than the weight of armanty meteorite located in Mongolia. If the sum of their weight is 88 tons. Find the weight of each.
Answer by Alan3354(69443)

The largest meteorite in the world is hoba west located in Namibia. It's [sic] weight is three times more than the weight of armanty meteorite located in Mongolia. If the sum of their weight is 88 tons. Find the weight of each.
-----------
h = weight of Hoba
a = weight of armanty
---
h = a + 3a = 4a
h + a = 88
4a + a = 88
5a = 88
a = 17.6 tons
---
h = 4a
h = 70.4 tons
---------
PS it's = it is